The most recent financial statements for GPS, Inc., are shown here: Income Statement Balance Sheet $23,600 Assets $119,000 Debt Sales Costs Taxable income Taxes (35%) $31,600 87,400 $119,000 15,800 $7,800 Total Equity $119,000 Total 2,730 $5,070 Net income Assets and costs are proportional to sales. Debt and equity are not. A dividend of $1,560 was paid, and the company wishes to maintain a constant payout ratio. Next years sales are projected to be $28,800. Required: What is the external financing needed?
Answer #1

EFN = (A/S) x (Δ Sales) - (L/S) x (Δ Sales) - (PM x FS x (1-d))

A/S = assets/sales = 119,000/23,600 = 5.04

L/S = liabilities/sales = 31,600/23,600 = 1.34

Change in sales (Δ Sales) = 28,800-23,600 = $5,200

PM = 5070/23600 = 0.21. FS = 28,800 and 1-d = 1-(1560/5070) = 0.69

Thus EFN = 5.04*5200 - 1.34*5200 - (0.21*28,800*0.69)

= 26,220.34 - 6,962.71 - 4,283.39

= 14,974.24